An ideal turbojet engine is analyzed using the cold air standard method. Given specific operating conditions determine the temperature, pressure, and enthalpy at each state, and the exit velocity.

Note: The specific heat ratio and gas constant for air are given as k=1.4 and R=0.287 kJ/kg-K respectively.

--Given Values--

T1 (K) = 236

P1 (kPa) = 56

V1 (m/s) = 217

rp = 11.1

rc = 1.78

a) Determine the temperature (K) at state 2.

b) Determine the pressure (kPa) at state 2.

c) Determine the specific enthalpy (kJ/kg) at state 2.

d) Determine the temperature (K) at state 3.

e) Determine the pressure (kPa) at state 3.

f) Determine the specific enthalpy (kJ/kg) at state 3.

g) Determine the temperature (K) at state 4.

h) Determine the pressure (kPa) at state 4.

i) Determine the specific enthalpy (kJ/kg) at state 4.

j) Determine the temperature (K) at state 5.

k) Determine the pressure (kPa) at state 5.

l) Determine the specific enthalpy (kJ/kg) at state 5.

m) Determine the temperature (K) at state 6.

n) Determine the specific enthalpy (kJ/kg) at state 6.

o) Determine the velocity (m/s) at the exit of the turbojet engine.
Answer #1

* Solution - PAGE at Considering the conditions with parameters from question. the He se T= 23.6% Pi 56 Epe Via Ritms OP 1!!!1s PAGE 2 251 XO 3.5 P2 = 56 236 = 56*(1.0975). -82-77 77.52X20 spefic enthalpy h2 = Co Tz (259) - 1005 ..42 = 260-295 65/13PAGE-3 State -4 8-1 Volume out To volumead $ 3 =) = 895.K B = ? - 86o a2 KP hy = 899 kJ/kg State- work done ont Work done onPAGE 4 try*** ho= hs-with (2129² = 646- 2 ho 241.90 og Vos Valp (TS-To) v (roos)(61272409:56) 68.856 mis VG -- ل PLEASE UPROT